OpenStudy (anonymous):

Find the zeros of the polynomial function and state the multiplicity of each. f(x) = 4(x + 7)2(x - 7)3

OpenStudy (aum):

f(x) will be zero when either of the factors is zero. x+7 = 0 or (x-7) = 0 solve for x.

OpenStudy (aum):

\[ f(x) = 4(x+7)^2(x-7)^3 \\ (x+7)^2 = 0 \text{ or } (x-7)^3 = 0 \\ x = -7 \text{ or } x = 7 \text{ are the roots.} \\ \text{Since } (x+7)^2 = 0, x = -7 \text{ has a multiplicity of 2} \\ \text{Since } (x-7)^3 = 0, x = 7 \text{ has a multiplicity of 3} \]

OpenStudy (anonymous):

Thank you!!!!